# Depth-Weighted Detection of Behaviours of Risk in People with Dementia using Cameras

### v1

@@ 5 Conclusion and Future Work
Caring for individuals with dementia in care settings is increasingly challenging due to growing patient numbers and understaffing issues. The behavioural and psychological symptoms of dementia can lead to incidents that jeopardize the health and safety of patients, staff, and caregivers. Leveraging existing video infrastructure can enable the development of innovative deep-learning approaches to detect such behaviours, thereby preventing injuries and enhancing patient care. As a step in this direction, in this paper, we proposed the depCAE method that uses a novel depth-weighted loss to enforce equivalent importance to both near and far pixels. The depCAE achieved AUROC of $0.852$, $0.81$ and $0.768$ for the three cameras, outperforming any existing methods and recorded lower FPR than baseline 3DCAE method for all three cameras, thus reducing the false alarms for detecting behaviours of risk in PwD. Further, we used the annotation of the training set outliers to determine the operating threshold. Both aspects were critical to depCAE performance and were validated through ablation analysis. This motivates further multidisciplinary research for the deployment of video surveillance-based behaviours of risk detection systems in nursing and care settings for PwD. Similar multidisciplinary approaches have shown to be helpful in medical settings [25, 26]. A video-based behaviours of risk detection system will help to improve the quality of life for the residents in care homes.
Future work can involve validating our approach on data collected from an LTC that typically encounters fewer behaviours of risk incidents. Pose tracking cameras that directly extract body joints information without recording actual scene [27] can be employed in private areas to detect behaviours of risk in a privacy-protecting manner [4, 28]. It can be a time-consuming and expensive task to train a method separately on each new scene or environment. A lot of time and resources can be saved if a method trained on one scene can generalize well in detecting behaviours of risk in an entirely new scene. Few-shot learning and transfer learning can be employed to improve the performance of depCAE for cross-camera generalization.

### v2

@@ 5 Conclusion and Future Work
Caring for individuals with dementia in care settings is increasingly challenging due to growing patient numbers and understaffing issues. The behavioural and psychological symptoms of dementia can lead to incidents that jeopardize the health and safety of patients, staff, and caregivers. Leveraging existing video infrastructure can enable the development of innovative deep-learning approaches to detect such behaviours, thereby preventing injuries and enhancing patient care. As a step in this direction, we proposed an approach that uses a novel depth-weighted loss to enforce equivalent importance to both near and far events and uses training outliers to determine the anomaly threshold. The depCAE achieved AUROC of $0.852$, $0.81$ and $0.768$ for the three cameras, outperforming other methods. The proposed approach recorded lower FPR than the existing methods across all three cameras, thus reducing the false alarms in detecting behaviours of risk in PwD. Further, the proposed approach consistently demonstrated its effectiveness and can easily be integrated into any existing autoencoder-based anomaly detection method. This motivates further multidisciplinary research for the deployment of video surveillance-based behaviours of risk detection systems in nursing and care settings for PwD. Similar multidisciplinary approaches have shown to be helpful in medical settings [31, 32]. A video-based behaviours of risk detection system will help to improve the quality of life for the residents in care homes.
Future work can involve validating our approach on data collected from an LTC that typically encounters fewer behaviours of risk incidents. The influence of camera placement on the global system performance needs to be further investigated by evaluating how the position, angle, and orientation of the cameras impact the system’s ability to accurately and reliably detect behaviours of risk. Further, the effect of utilizing the depth map as part of data as an additional input channel can be investigated. 3D object reconstruction-based approaches [33, 34] can be used to circumvent the problem of occlusions and improve the performance of detecting behaviours of risk as anomalies. Pose tracking cameras that directly extract body joints information without recording actual scene [35] can be employed in private areas to detect behaviours of risk in a privacy-protecting manner [4, 36]. It can be a time-consuming and expensive task to train a method separately on each new scene or environment. A lot of time and resources can be saved if a method trained on one scene can generalize well in detecting behaviours of risk in an entirely new scene. Fine-tuning on target camera using few-shot learning and transfer learning can be employed to improve the performance of depCAE for cross-camera generalization.